Mary Kim invested $964 in 18 shares of stock. A month
later, she sold the 18 shares at $61.48 per share. She also
invested in 350 shares of stock for a total of $27,422.50.
She sold this stock for $85.35 a share and paid $495 in
total commissions. What was Mary Kim’s profit or loss
on the transactions to the nearest dollar?

Answer by ewatrrr(24785)


Hi,
Profit earned for each transaction
18*61.48 - 964 = 142.64 total profit
350*85.35 - 27422.50= 2450 total profit
Total profit:
2450.00 + 142.64 - 495.00 = $2097.64